Suppose that an amount in dollars, is invested in a private financial institution, with interest

compounded continuously at 8% per year.

a) Write the equation in terms of P, and 0.08 where P, is the starting amount

invested. And the final balance in the account is denoted with variable P

b) Suppose that $2000 is invested. What is the total amount in the account after 3

years?

c) How many years will it take to have more than the invested amount



1
Expert's answer
2021-06-16T04:21:20-0400

Continuous compoynding is given by the equation:

"FV=P_oe^{rn}"

Where FV = future or final value

"P_o=" present or initial value invested

e= rate of increase

r = interest rate

n = number of years


a)"FV=P_oe^{rn}"

e= 2.71828, n= 1, "P_0=FV=P" , r= 0.08

"\therefore\frac{P}{P_o}=2.71828^{0.08×1}"

"\frac{P}{P_o}=2.71828^{0.08}"

b) Assume that P =$2000 and n= 3


"FV_3=2000(2.71828)^{0.08\u00d73}"

"FV_3=\$2542.50"


c) From (a), "\frac{P}{P_o}=e^{rn}"

"P>P_o" at "n+1"